# Multi-Step Reasoning with Large Language Models, a Survey

### v2

@@ 4.3.2. What Can LLMs Not Do?
Now that grade school math word problems are largely solvable, harder reasoning benchmarks in other domains are appearing (Ahn et al., 2024). Most of the reasoning capabilities exhibited by LLMs are due to the representational powers of the transformer architecture, and how in-context learning is able to harness them. Prompt engineering and prompt control play a crucial role in the kind of reasoning that we have seen in the papers. Models can be instructed to write their own reasoning prompts; however, such Auto-GPT or Auto-CoT prompts need careful evaluation, verification, and grounding in the real world, to prevent degeneration into a hallucinatory world of their own. Models can also be instructed to interact with the world, and become the tool of external scaffolding that evaluates, controls and improves the prompts (Plaat et al., 2025). Some of what we experience as reasoning by the LLM, is controlled by the prompt or the scaffolding algorithm. Studies into in-context reinforcement learning aim to answer the question if prompt learning is able get the LLM to create a prompt to exhibit dynamic reasoning by itself (Lee et al., 2023; Schultz et al., 2024; Demircan et al., 2024; Paglieri et al., 2024).
@@
Other works study the dangers of the size of LLMs. Bender et al. (2021) mention the environmental risks associated with the large computational training demands, as well as the difficulty of understanding the training data, for example in the context of bias. Furthermore, there are ethical, legal, and copyright concerns regarding the data that LLMs are trained on. Finally, to prevent putting too much trust in the outcome of LLMs, we should understand their failure modes better, such as the well-publicized problems of hallucination (inventing facts that look right but are not). Here, mechanistic interpretability can be used to explore LLM representations and understand where they go wrong (Bereska and Gavves, 2024; Rai et al., 2024; Chen, 2025).
@@ 4.3.3. Reasoning towards Metacognition
Many reasoning approaches highlight self-reflective aspects in their names and in how they work. The prompts that prompt the models to reason are being improved with the outcome of the reasoning process, and in Buffer-of-thoughts thought-templates are used that are derived from other reasoning processes. Wang and Zhao (2024) study Metacognitive-prompting. Inspired by Chain-of-thought and Self-consistency, they create manually designed prompts to increase the understanding of language models. Another work, again inspired by Chain-of-thought and Self-consistency, connects psychology and LLMs. Didolkar et al. (2024) study metacognitive capabilities of LLMs in mathematical problem solving, both on GSM8K and on the harder MATH problems (Hendrycks et al., 2021). First, the model is prompted to find a skill name for each problem instance in the dataset. For 7000 instances of GSM8K, 500 skill names were found by the model. Next, these 500 names are clustered down to 22 skills. They find that by using the names of these 22 skills in Chain-of-thought-like prompts, more problems are solved than with standard Chain-of-Thought/Self-consistency/PAL prompts. Examples of the 22 skill names are multiplication-and-addition, basic-arithmetic, subtraction, and algebra. Interestingly, the authors find that the skill exemplar repository that is trained on a strong model (GPT-4), also down-translates to a weak model (GPT-3). The performance of the weaker model benefits from the skill-name-enhanced prompts. Begus et al. (2025) propose a program for the study of metalinguistic abilities of LLMs. LLMs struggle with analogies (Lewis and Mitchell, 2024; Mitchell and Hofstadter, 1990) and analogical reasoning. Metacognitive reasoning with LLMs is still in its early stages.
@@ 4.4. LLMs that Generate Computer code
LLMs can also be used in algorithmic mode, or Chain-of-thought mode, where an extra element is added to the prompt that gives an example of which steps to take to arrive at the answer (Sel et al., 2023). Lee et al. (2023); Wang et al. (2024c); Schultz et al. (2024) show that the transformer architecture can be trained such that search and reinforcement learning algorithms can be executed in-context. (LLMs pretrained on natural language are typically not good at in-context reinforcement learning (Ruoss et al., 2024; Paglieri et al., 2024)).
We have seen that LLMs are not good at preserving state between calls (Kempinski et al., 2025). External algorithms can be added to manage the state, that call the LLM in an external mode (Yao et al., 2024, 2023). Tree-of-thoughts, and similar approaches, use the LLM in a stateless functional way, with the prompt containing the instructions, not unlike a natural language variant of a functional programming language such as LISP or Haskell.
Finally, an LLM can be used in code generation mode. The prompt specifies a problem, and the LLM is asked to generate computer code (such as Python or PDDL). The code may contain variables that manage the state correctly. The code can be executed directly, or first be checked and improved by a programmer. We expect that the use LLMs as code generators, and prompt-engineering as natural-language-programming, will continue to grow.
@@ 4.5. Research Agenda
  • In Context Reinforcement Learning—Search control beyond greedy search is often implemented as an external reinforcement learning algorithm. Is it possible to incorporate the control stage of the reasoning pipeline into one static prompt, as in-context reinforcement learning (ICRL)? Studies indicate that models can be trained for decision making (Lee et al., 2023; Demircan et al., 2024), but general LLMs have not yet been shown to perform ICRL well (Schultz et al., 2024; Paglieri et al., 2024; Ruoss et al., 2024; van Wetten et al., 2025).
@@
  • Faithfulness—Our theoretical understanding of prompt-based reasoning with LLMs is incomplete. The research on faithfulness highlights one example of our lack of understanding. In general, more insight into the working of multi-step in-context learning in LLMs is dearly needed.
  • Smaller language models—Efficiency is an important element for wide adoption of language models. Smaller models have many advantages over larger models: a smaller environmental footprint, more accessible to people and organizations with fewer computational resources, and the possibility to finetune them. Models may have fewer parameters, or the parameters may be quantized with fewer bits. Unfortunately, with a smaller number of parameters, the models also have less reasoning power. It is important for future work to investigate this trade-off. Solutions to compensate the limited capabilities of smaller models are distillation of reasoning to small language models and reinforcement learning approaches (Shao et al., 2024) to improve the efficiency of the finetuning/reasoning pipeline (Guo et al., 2025; Du et al., 2025).
@@ 5. Conclusion
Prompt-based in-context learning is an efficient machine learning method, requiring no parameter updates to the LLM. While achieving good performance on language tasks, performance on reasoning tasks was lacking originally. Reasoning tasks, such as math word problems, are typically solved in a step-by-step fashion. Recently prompts have been developed that guide an LLM to “think step by step” (Chain-of-thought), and to evaluate and verify the step results. The performance of multi-step reasoning with LLMs has improved greatly, and the field has progressed beyond math word problems. Together, the surveyed methods allow the LLM to follow high-quality multi-step reasoning chains. Python code or other formal languages have been used successfully to reduce the error in reasoning steps. Also, in the field of autonomous agents and robotic action, good performance has been achieved by grounding reasoning answers in the environment and the physical constraints of robotic movement.
For complex reasoning tasks a large number of reasoning steps may be generated. To control the size of the multi-step reasoning space dynamically, external scaffolding algorithms can be used. Often, variations on search algorithms are used, and especially reinforcement learning. The symbolic and connectionist AI traditions come together in reasoning prompts and search algorithms that help LLM neural networks solve natural language math word and related problems. Inference-time reasoning results can be used to augment finetuning, in a feedback loop.
LLMs hallucinate and suffer from bias, and their use poses different ethical dangers. In multi-step reasoning methods, self-verification methods have been developed to reduce error-accumulation, yet ethical dangers may remain, especially since reflection is not error-free. Additionally, retrieval augmentation methods ground LLM output directly in sources such as Wikipedia.
Among the most popular reasoning benchmarks in this survey is GSM8K, which contains 8500 grade school math word problems. With older LLMs such as GPT-3, reasoning approaches show an improvement of 20-50% points over standard prompting methods, and some even more. Many current LLMs have since adopted the reasoning approaches. The success of reasoning with LLMs has attracted more applications, and with them, benchmarks are diverging.
The field of reasoning with LLMs is quite new, and theoretical understanding is lacking in important areas, such as faithful reasoning (models may sometimes find the right answer using an incorrect reasoning chain). There may be more opportunities for efficiency gains by integrating the training of reasoning models, using reinforcement learning, both in training and in context. Although prompt-based learning allows few-shot learning at inference time, the computational needs of LLM pretraining and finetuning are still high, hence the interest in small language models. Reasoning skills that work in large models can often be distilled to small models.
@@
Human thought is capable of metacognition, we can think about our thinking process. Many of the names of the approaches in this survey suggest a link to metacognition (Reflexion, Self-refine, Self-improvement, Inner-monologue). The first preliminary experiments of language models that reason about their reasoning skills have appeared.
@@
LLM-reasoning is an active field of research that shows great progress. Based on current limitations and open questions we provide a research agenda highlighting opportunities for further progress with reinforcement learning (for finetuning, in context reinforcement learning and self-reflection), for LLMs that generate code for external tools, for multimodal world models, for value-aligned reasoning, and for small language models, among others.
@@ Acknowledgements
We thank the anonymous reviewers for their valuable suggestions that have considerably improved the paper.

### v3

@@ 4.3.2. What Can LLMs Not Do?
Now that grade school math word problems are largely solvable, harder reasoning benchmarks in other domains are appearing (Ahn et al., 2024). Most of the reasoning capabilities exhibited by LLMs are due to the representational powers of the transformer architecture and how in-context learning is able to harness them. Prompt engineering and prompt control play a crucial role in the reasoning that we have seen in this survey. Models can be instructed to write their own reasoning prompts; however, such Auto-GPT or Auto-CoT prompts are prone to feedback-loop problems, and need careful evaluation, verification, and grounding in the real world, to prevent degeneration into a hallucinatory world of their own. Models can also be instructed to interact with the world, and become the tool of external scaffolding that evaluates, controls and improves the prompts (Plaat et al., 2025). Some of what we experience as reasoning by the LLM, is controlled by the prompt or the scaffolding algorithm. Studies into in-context reinforcement learning aim to answer the question if prompt learning is able get the LLM to create a prompt to exhibit dynamic reasoning by itself (Lee et al., 2023; Schultz et al., 2024; Demircan et al., 2024; Paglieri et al., 2024).
@@
Other works study the dangers of the size of LLMs. Bender et al. (2021) mention the environmental risks associated with the large computational training demands, as well as the difficulty of understanding the training data, for example in the context of bias. Furthermore, there are ethical, legal, and copyright concerns regarding the data that LLMs are trained on. Finally, to prevent putting too much trust in the outcome of LLMs, we should understand their failure modes better, such as the well-publicized problems of hallucination (inventing facts that look right but are not). Here, mechanistic interpretability can be used to explore LLM representations and understand where they go wrong (Bereska and Gavves, 2024; Rai et al., 2024; Chen et al., 2025).
@@ 4.3.3. Reasoning towards Metacognition
Many reasoning approaches highlight self-reflective aspects in their names and in how they work. The prompts that prompt the models to reason are being improved with the outcome of the reasoning process, and in Buffer-of-thoughts thought-templates are used that are derived from other reasoning processes. Wang and Zhao (2024) study Metacognitive-prompting. Inspired by Chain-of-thought and Self-consistency, they create manually designed prompts to increase the understanding of language models. Another work, again inspired by Chain-of-thought and Self-consistency, connects psychology and LLMs. Didolkar et al. (2024) study metacognitive capabilities of LLMs in mathematical problem solving, both on GSM8K and on the harder MATH problems (Hendrycks et al., 2021). First, the model is prompted to find a skill name for each problem instance in the dataset. For 7000 instances of GSM8K, 500 skill names were found by the model. Next, these 500 names are clustered down to 22 skills. They find that by using the names of these 22 skills in Chain-of-thought-like prompts, more problems are solved than with standard Chain-of-Thought/Self-consistency/Program-aided-language prompts. Examples of the 22 skill names are multiplication-and-addition, basic-arithmetic, subtraction, and algebra. Interestingly, the authors find that the skill exemplar repository that is trained on a strong model (GPT-4), also down-translates to a weak model (GPT-3). The performance of the weaker model benefits from the skill-name-enhanced prompts. Begus et al. (2025) propose a program for the study of metalinguistic abilities of LLMs. LLMs struggle with analogies (Lewis and Mitchell, 2024; Mitchell and Hofstadter, 1990) and analogical reasoning. Metacognitive reasoning with LLMs is still in its early stages.
@@ 4.4. LLMs that Generate Computer code
LLMs can also be used in algorithmic mode, or Chain-of-thought mode, where an extra element is added to the prompt that gives an example of which steps to take to arrive at the answer (Sel et al., 2023). Muennighoff et al. (2025) show how such test-time scaling can trade-off model size and training time. Lee et al. (2023); Wang et al. (2024c); Schultz et al. (2024) show that the transformer architecture can be trained such that search and reinforcement learning algorithms can be executed in-context.
LLMs do not preserve state between calls, and in order to allow for a natural dialogue to occur, chatbots such as ChatGPT insert the most recent history of preceding calls into each prompt before the model is called (Brown et al., 2020). For harder strategic reasoning tasks more advanced state management solutions are necessary, that call the LLM in an explicit external mode (Yao et al., 2024, 2023). Approaches such as Tree-of-thoughts use the LLM in a stateless functional way, with the prompt containing the instructions, not unlike a natural language variant of a functional programming language such as LISP or Haskell.
Finally, an LLM can be used in code generation mode. The prompt specifies a problem, and the LLM is asked to generate computer code (such as Python or PDDL), to circumvent the knowing-doing gap (Paglieri et al., 2024). The code may contain variables that manage the state correctly. The code can be executed directly, or first be checked and improved by a programmer. We expect that the use LLMs as code generators, and prompt-engineering as natural-language-programming, will continue to grow.
@@ 4.5. Research Agenda
Specific research topics for reasoning with LLMs are:
  • In Context Reinforcement Learning—Search control beyond greedy search is often implemented as an external reinforcement learning algorithm. Is it possible to incorporate the control stage of the reasoning pipeline into one static prompt, for implicit reasoning or in-context reinforcement learning (ICRL)? Studies indicate that models can be trained for decision making (Lee et al., 2023; Demircan et al., 2024), but general LLMs struggle to perform ICRL well (Schultz et al., 2024; Paglieri et al., 2024; Ruoss et al., 2024; van Wetten et al., 2025).
@@
  • Faithfulness—Our theoretical understanding of prompt-based reasoning with LLMs is incomplete. The research on faithfulness highlights one example of our lack of understanding. In general, more insight into implicit reasoning (Li et al., 2025) and the working of multi-step in-context learning in LLMs is dearly needed.
  • Smaller language models—Efficiency is an important element for wide adoption of language models. Smaller models have many advantages over larger models: a smaller environmental footprint, more accessible to people with fewer computational resources, and the possibility to finetune them. Models may have fewer parameters, or the parameters may be quantized with fewer bits. Unfortunately, with a smaller number of parameters, the models also have less reasoning power. It is important for future work to investigate this trade-off.
@@ 5. Conclusion
When large language models of sufficient size are prompted with examples, they can perform few-shot learning in-context, providing an answer at inference time, without retraining model parameters. Although they achieve good performance on language tasks, simple prompting methods do not perform well on reasoning tasks—tasks that humans typically solve in a step by step fashion. Chain-of-thought is an in-context prompting method to guide an LLM to “think step by step.” Chain-of-thought was originally developed for solving grade school math word problems. GSM8K, the most popular reasoning benchmarks in this survey, contains 8500 grade school math word problems. With older LLMs such as GPT-3, reasoning approaches show an improvement of 20-50% points over standard prompting methods. This success has spawned many new reasoning approaches. A potential problem in the Chain-of-thought method is that errors may accumulate over multiple reasoning steps. Prompting the LLM to reformulate problems in Python code (or another formal language) can successfully reduce the error rate. The success of reasoning methods has attracted more applications, and with them, benchmarks are diverging. In the field of autonomous agents and robotic action, good performance has been achieved by grounding reasoning answers in the environment and the physical constraints of robotic movement.
Many current LLMs have adopted Chain-of-thought approaches. In this survey we categorize the approaches on how they generate, evaluate, and control the reasoning steps. Inference-time reasoning methods are also used for LLM finetuning. Reinforcement learning with verifiable rewards (RLVR) and group relative policy optimization (GRPO) use inference-time reasoning results to augment finetuning for reasoning tasks.
@@
For complex tasks the number of reasoning steps that is generated may be large. The size of the multi-step reasoning space can be controlled dynamically by external search or reinforcement learning algorithms, often in the form of a wrapper of Python code that generates LLM-prompts. Many of the names of the approaches in this survey suggest a link to metacognition (Reflexion, Self-refine, Self-improvement, Inner-monologue)—the act of thinking about one’s own thought process. The first preliminary experiments of language models that reason about their reasoning skills have appeared.
The field of reasoning with LLMs is quite new, and theoretical understanding is lacking in important areas, such as faithful reasoning (models may sometimes find the right answer using an incorrect reasoning chain). LLMs hallucinate and suffer from bias, and their use poses ethical and societal dangers. Self-verification methods have been developed to reduce error-accumulation, and retrieval augmentation methods (RAG) ground LLM output directly in sources such as Wikipedia. However, ethical and societal dangers remain, especially since also reflection is not error-free.
Although prompt-based learning allows few-shot learning at inference time, the computational needs of LLM pretraining and finetuning are still high, hence the interest in small language models. Reasoning skills that work in large models can often be distilled to small models.
@@
LLM-reasoning is an active field of research that shows great progress. Based on current limitations and open questions we provide a research agenda highlighting opportunities for further progress. Among the items mentioned are reinforcement learning for finetuning and self-reflection, LLM agents that generate code for external tools, multimodal world models, value-aligned reasoning, and small language models.
@@ Acknowledgements
We thank the anonymous reviewers for their valuable suggestions that have considerably improved the article.